### v1

# AI-Powered Episodic Future Thinking

Abstract
Episodic Future Thinking (EFT) is an intervention that involves vividly imagining personal future events and experiences in detail. It has shown promise as an intervention to reduce delay discounting—the tendency to devalue delayed rewards in favor of immediate gratification—and to promote behavior change in a range of maladaptive health behaviors. We present EFTeacher, an AI chatbot powered by the GPT-4-Turbo large language model, designed to generate EFT cues for users with lifestyle-related conditions. To evaluate the chatbot, we conducted a user study that included usability assessments and user evaluations based on content characteristics questionnaires, followed by semi-structured interviews. The study provides qualitative insights into participants’ experiences and interactions with the chatbot and its usability. Our findings highlight the potential application of AI chatbots based on Large Language Models (LLMs) in EFT interventions, and offer design guidelines for future behavior-oriented applications.
1 Introduction
Approximately 40% of adults in the United States are currently living with obesity, with projections suggesting that this figure will reach 50% by 2030 [1, 2]. Although some progress has been made in addressing obesity, many communities still lack access to prevention and treatment resources, worsening health disparities related to obesity [3]. Losing weight successfully depends on making choices today that are influenced by long-term goals, as most of the health benefits from weight loss come after months or even years of maintaining consistent healthy habits.
Episodic future thinking (EFT) involves vividly imagining personal future events and experiences in detail [4]. EFT has been increasingly used as an intervention for behavior change in various maladaptive health behaviors and populations with lifestyle-related conditions to reduce delay discounting [5]. Delay discounting (DD) refers to the tendency to choose immediate reward over greater, future benefits [6]. EFT interventions usually involve participants identifying several positive future events they look forward to and plan to attend at different times. They write or describe these events in detail and are encouraged to vividly imagine them when making decisions involving future trade-offs [7]. Personalized positive future cues have been shown to significantly impact decision-making by encouraging individuals to consider longer-term consequences. These cues help expand the "temporal window," which refers to the ability to think further into the future when evaluating choices [8, 9, 10].
Research indicates that obesity is linked to high levels of DD. For instance, high levels of DD are both currently and over time linked to having a higher body mass index (BMI), consuming more calories, and engaging in less physical activity [11, 12, 13]. Additionally, experimental studies demonstrate that interventions aimed at altering DD lead to related improvements in dietary habits, further supporting the idea of a functional connection between the two [13, 14, 15, 16].
Laboratory studies have shown that engaging in EFT for a short time can effectively reduce DD in adults with obesity [17, 18, 19, 20, 21], and a range of other clinical disorders [22] . This means EFT helps people prioritize long-term goals over immediate rewards. Studies also demonstrate that brief EFT sessions can decrease the purchase and consumption of unhealthy, high-calorie foods both in controlled lab settings and in real-life situations, such as grocery shopping or dining in public [14, 23, 16, 24].
Building on these findings, EFT has been adapted into a remotely delivered intervention as part of clinical trials for managing type 2 diabetes. This intervention aims to help individuals adhere to lifestyle changes recommended for treating obesity and its comorbidities [25, 26, 5]. EFT works by helping patients stay focused on their long-term health goals while reducing the appeal of immediate temptations like unhealthy foods or inactivity, which could derail their progress.
There are two common approaches to creating EFT cues for studies on delay discounting and health behaviors: interview-guided and survey-guided methods [17, 27]. Both methods share some similarities in how the cues are generated. For instance, researchers may give specific instructions based on the study’s goals and participants’ behavior [5]. In clinical practice, having clinicians conduct interviews to guide patients through EFT is costly, requires significant training to ensure treatment fidelity, and limits when and where patients can access the intervention. While self-administered, survey-based approaches address these challenges by reducing costs and increasing accessibility, they introduce new concerns about maintaining consistency and effectiveness in treatment (e.g., due to terse or off-topic cues). To address these issues, this study aims to evaluate an artificial intelligence (AI)-based method for delivering EFT. This approach would improve consistency, scalability, accessibility, and effectiveness, allowing EFT to be implemented on a much larger scale while minimizing costs and other barriers. We aim to enhance and evaluate an existing AI-based chatbot [28], referred to as EFTeacher, for EFT intervention. EFTeacher is a chatbot built on GPT-4-turbo, a Large Language Model (LLM). EFTeacher is designed to assist participants in identifying personally meaningful future events, and in creating for them detailed text descriptions, or cues, for those events.
In this study, we seek to answer the following research questions:
  • RQ1: What are users’ overall experiences of using EFTeacher?
  • RQ2: What are users impressions of using EFTeacher?
  • RQ3: How do users perceive working with an interactive AI chatbot for EFT generation?
To address RQ1 and RQ2, we adopt a quantitative approach utilizing the System Usability Scale (SUS) [29], and an analysis of specific content characteristics [30]. We answer RQ3 by using thematic analysis [31, 32].
2 Related Work
Conversational agents have emerged as valuable tools for delivering cost-effective and practical behavioral interventions. Studies demonstrate their effectiveness in promoting physical activity, encouraging healthy diets [33, 34], and supporting weight loss efforts [35]. The design of engaging and impactful chatbot interventions remains a crucial focus for maximizing their potential [36]. Furthermore, AI-driven conversational agents have been used in behavior change systems to create more interactive and engaging interventions [37, 38, 39, 40].
Human-computer interaction (HCI) technologies have shown promise in fostering mindfulness and enhancing well-being [41]. Studies highlight its ability to increase relaxation, focus, and self-awareness, thereby improving decision-making and motivating behavior change [41]. Mindful interaction with technology emphasizes the importance of reflection and deliberate choices during digital interactions, further enhancing personal growth and well-being. As a result, researchers have explored various approaches to designing conversational agents tailored to specific behavioral goals.
Physical Activity and Healthy Eating: Olafsson et al. [42] compared the impact of humorous versus non-humorous virtual agents in promoting exercise and healthy eating, finding humor to enhance motivation and engagement significantly. Similarly, Sun et al. [43] investigated the impact of humor in chatbot-delivered interventions designed to increase physical activity. The study focused on user interactions with a chatbot, analyzing levels of participation, motivation, and physical activity. The findings showed that the use of humor in the chatbot helped users become more physically active. Paay et al. [44] highlighted the role of digital personal assistants (DPAs) in motivating physical activity such as exercise, with virtual rewards being particularly effective. Kocielnik et al. [45] introduced a reflection tool called the “Reflection Companion,” which encourages users to notice patterns in their physical activity and plan future actions. The system uses short daily conversations and visual data representations to help users recognize their habits and strategize for improvement, underscoring the value of reflective practices in promoting healthy behaviors.
Smoking Cessation: Alphonse et al. [46] revealed that human-like traits in chatbots promote accountability and engagement, making them effective tools for smoking cessation. Their qualitative study explored smokers’ experiences with Quit Coach, a quick-response chatbot integrated into a smoking cessation app. The findings suggest that the chatbot’s design increased users’ sense of accountability and engagement, supporting the use of chatbots in smoking cessation interventions. Kjaerulff et al. [47] explored voice-based user interfaces (VUIs) that facilitated mindfulness techniques for smoking cessation. Participants used the VUI to track cravings, engage in mindfulness exercises, or reflect on smoking habits by listening to pre-recorded audio responses. Results indicated that both VUIs and augmented VUIS (AVUIs) show promise in supporting smoking cessation and proposed integrating Large Language Models (LLMs) for enhanced personalization and adaptability.
Applications in Chronic Disease Management: Conversational agents have also shown promise in managing chronic conditions like diabetes. Bin et al. [48] proposed a chatbot design framework that addresses limitations in existing systems. This framework was developed using surveys, interviews, and a review of current healthcare chatbots. By incorporating insights from the interviews, the framework was designed to meet user needs and preferences, thereby improving the chatbot’s effectiveness in preventing type 2 diabetes. Baptista [49] evaluated the acceptability of an embodied conversational agent (ECA) for type 2 diabetes self-management. Using both survey data and participant interviews, the study found that users appreciated the agent’s friendly and motivational support, though improvements in communication and appearance were recommended, highlighting the potential of ECAs in chronic disease management.
Supporting Well-Being and Mindfulness in Food-Related Activities: Studies have also examined how conversational agents can enhance well-being during food-related activities. Parra et al. [50] explored the potential of agents to promote mindfulness during meals, especially following the impact of COVID-19 on eating habits. The study began by examining how the COVID-19 pandemic affected young adults’ eating habits, noting its negative impact on well-being. Based on these findings, the researchers developed two interactive systems. Both systems demonstrated that conversational agents could make these activities more enjoyable and foster mindfulness. The researchers aim to develop more advanced agents using LLMs and activity recognition to adapt to users’ needs effectively. Zhang et al. [51] presented a pilot study investigating how voice assistants can encourage mindful eating. Researchers created two versions of a voice assistant app: one with a friendly approach and another with a counseling approach. They tested these apps to evaluate user preferences and how the assistant’s social role influenced their mindful eating experience. The study found that participants preferred the "friend" version, offering insights for designing voice assistants that support mindfulness practices.
Advancing Conversational Agents for Personal Growth: Several studies focus on refining conversational agents to support broader personal development goals. Saravanan et al. [52] explored the design and evaluation of a conversational system for a social robot employing motivational interviewing to encourage healthy eating. Their findings demonstrated that referencing past interactions in motivational interviewing sessions improved user engagement and behavior change. Participants interacted with the virtual robot over three sessions, during which their motivation levels and goal progress were measured. Results indicated that varied references to previous sessions significantly increased motivation for behavior change. Meyer et al. [53] highlighted the potential of conversational agents trained to recognize behavior-change language using motivational interview concepts. Browne et al. [54] developed and tested a conversational AI health coaching system integrated into an app. The study evaluated the system’s performance, user interaction, and the effectiveness of prompt engineering techniques. They assessed the app’s impact on mental well-being, finding improvements in positive affect, reduced negative affect, and enhanced self-reflection. The study concludes with a discussion of the app’s usability, user feedback, and suggestions for future improvements.
Large Language Models in Conversational Systems: Recent advancements in LLMs have further expanded the capabilities of conversational agents. Arakawa et al. [55] explored LLM-powered chatbots in executive coaching, emphasizing their accessibility and logical reasoning while highlighting the importance of human-AI collaboration to promote deep self-reflection. James et al. [56] evaluated LLMs for generating personalized goals in gamified mHealth apps. A study comparing LLM-generated and pre-determined goals highlighted the scalability and efficiency of LLM-driven content creation. Nepal et al. [57] introduced the MindScape app, which uses AI-generated journaling prompts tailored to smartphone behavioral data. The study showed improvements in mental well-being and self-reflection, emphasizing the app’s potential for enhancing traditional journaling for effective interventions that promote self-awareness and holistic well-being.
Our study aims to perform a qualitative evaluation of EFTeacher, which is based on a Large Language Model (LLM) for EFT cue generation. To the best of our knowledge, we are the first group to leverage LLMs for EFT cue generation, which has potential applications in interventions addressing maladaptive health behaviors.
3 Study Method
We developed an AI chatbot [28], EFTeacher, which is modeled after the clinician-led EFT interviews [17, 16, 20] and self-administered surveys [19, 18, 58] used previously. We implemented a website [59] for a single-session study in which participants were given a username and password and could log into the website as shown in Figure 1 to complete the AI-assisted EFT interview. We used GPT-4-Turbo [60] through Microsoft Azure, which is a university-approved cloud service for EFTteacher. As shown in Figure 1, memory, in the form of a list, is the history of the conversation which is passed into the prompt along with the user query (message) so the chatbot can generate an output. As depicted in the diagram, the memory stores both the responses generated by the chatbot and the corresponding user queries. Each interaction between the human and GPT-4 is added to this list so that it can be used in the next input for the model.
Figure 1: User Study System
On the chat page of the website, depicted in Figure LABEL:wireframe, users interacted with the chatbot, which guided them in generating a set of EFT cue texts. EFTeacher first introduced EFT cues to participants. They were then provided with clear criteria for crafting a good EFT cue. Participants were guided to ensure their cues met specific criteria: the event should be something they look forward to, vividly imaginable, confined to a single 24-hour period, and written in the present tense as if they are currently experiencing it. To help them understand, participants were shown examples of well-crafted EFT cues that met all the criteria, as well as poor examples. They were then prompted to create their own EFT cues for specific timeframes, starting from one month into the future, ensuring their events were realistic, specific, and followed the given guidelines. This step-by-step approach ensured participants understood EFT cues and could create effective ones.
For each time-frame, participants first provided brief contextual statements about the event. Then they answered a series of standardized questions to add specific details, such as where the event was taking place, who was present, what actions were occurring, and how they felt in that moment. EFTeacher then combined these brief statements and detailed responses into a coherent cue, typically 3-4 sentences long. If participants gave incomplete or unrelated answers, EFTeacher prompted them to revise their responses to ensure the quality and relevance of the cues. Once an EFT cue was generated, EFTeacher sought feedback from the participant to adapt the cue based on their preferences, ensuring it aligned with their personal perspective and needs. The user and chatbot interacted iteratively until a desired EFT cue was generated and accepted by the user. Once finalized, the EFT cue was extracted through a predefined routine and displayed in a dedicated box (cue box), as shown in Figure LABEL:wireframe. As the conversation progressed and an EFT cue was generated for any time-frame, users could view the cues in the cue box, navigate back to previous cues, and review them. This allowed users to keep track of which events had already been discussed during the EFT generation For this study, once users completed covering events for seven different time-frames (one month, three months, one year, three years, five years, and ten years), the conversation concluded.
Upon completing the conversation, the chatbot’s logs, including the entire conversation (referred to as memory), were saved as a file in the database. A designated routine stored the extracted EFT cues in the database, and then sent them to Qualtrics. Following this, users were redirected to the next page, where they evaluated the system’s usability using the System Usability Scale (SUS) [29] questions. After completing these questions, they were redirected to the Qualtrics website. In Qualtrics, participants were presented with the generated EFT cue texts by EFTeacher, organized by the different time-frames. For each cue, participants were asked to evaluate its characteristics on a 5-point Likert scale, ranging from 1 (not at all) to 5 (extremely). They rated each cue based on four key aspects: Liking, Importance, Excitement, and Vividness. Afterward, the participants were presented with a survey to provide additional feedback about their experience by answering open-ended questions, as explained in detail in Section 3.2.
3.1 Participants
Following approval from the Institutional Review Board (IRB), the EFT research team carried out an evaluation study with 25 participants, specifically targeting individuals with obesity. Eligibility was assessed based on the following criteria: a BMI of 30 or higher, fluency in English, and an age range of 18 to 65 years. Participants were invited by email from a database of individuals who had previously completed eligibility questionnaires for studies at the center for our prior studies. All participants had provided consent to be contacted about research studies. The recruitment process began with 346 email invitations of which 17.9% (62 individuals) started the screening questionnaire—to evaluate eligibility. Out of those, 15.3% (53 individuals) completed the questionnaire, resulting in 9.8% (34 individuals) being deemed eligible and providing consent to participate. Ultimately, 7.2% (25 individuals) completed the study. The demographics of the participants are detailed in Table 1. Household size is summarized using the median and interquartile range (IQR), where the median represents the middle value and the IQR (25th–75th percentile) shows the range in which the central 50% of the data falls. In general, the set of participants seemed to be relatively balanced across key demographic characteristics.
Table 1: Summary of Participant Demographics
Characteristic Complete (n=25)
BMI (kg/m2), mean (SD) 38.8 (8.1)
Age, mean (SD) 38.5 (10.1)
Gender, n (%)
Male 10 (38.5)
Female 15 (61.5)
Household income, n (%)
<<<50,000 7 (30.1)
\$50,000–\$79,999 6 (23.1)
\$80,000–\$119,999 7 (27.0)
$\geq$120,000 5 (19.2)
Household size, median (IQR) 2 (2, 4)
Education, n (%)
HS/GED or less 7 (30.1)
Some college 6 (23.1)
2-yr college degree 3 (11.5)
4-yr college degree 7 (26.9)
Grad/professional degree 2 (7.7)
Race, n (%)
White 17 (65.4)
Black/AA 3 (15.4)
Asian 1 (3.8)
More than one race 4 (15.4)
Ethnicity, n (%)
Hispanic/Latino 2 (7.7)
Glycemic status, n (%)
Type 2 diabetes 11 (42.3)
Prediabetes 6 (23.1)
Normoglycemia 8 (34.6)
Users began by completing an informed consent process on the Qualtrics platform [61] where they provided their demographic information. They were then redirected to our website, where each user was assigned a unique ID stored in the database for use during their interactions.
3.2 Data Collection
In this study, one of the goals was to better understand people’s experiences with EFTeacher, particularly their interactions with it. To do this, we used thematic analysis to perform a qualitative analysis, identifying themes from open-ended survey responses. This approach was inductive, meaning themes were drawn directly from the participants’ responses without being predetermined.
Participants were asked five open-ended questions to share their thoughts and feelings in their own words about using EFTeacher. To add context to these responses, participants first rated their agreement with several statements on a scale from “Strongly Disagree” to “Strongly Agree.” These statements focused on how EFTeacher’s responses influenced their ability to imagine and describe future events. For example, prompts included questions like whether EFTeacher’s responses felt human-like, helped them imagine future events, or made those events more vivid.
After these ratings, participants were asked to elaborate on why they felt the way they did. They were also asked to identify features of EFTeacher they found helpful and engaging, as well as those they found frustrating or less useful. This included listing at least two features they liked and believed should remain in future versions of the chatbot and two features they disliked and thought should be improved. Finally, participants could share any additional feedback or thoughts about their overall experience with EFTeacher. This process aimed to gather detailed insights into the user experience, including both positive and negative aspects. This feedback would help refine EFTeacher and inform the design and evaluation of similar chatbots in the future. Each participant is assigned an ID number from 100-150.
Quantitative data was also collected for both SUS and content characteristics. For the SUS [29], we calculated the mean and standard deviation (SD) across all 25 users. For key content characteristics (i.e., Vividness, Liking, Excitement, and Importance), we calculated the median of these four items from each EFT cue that was generated for a specific timeframe, where users rated each item on a scale of 1 (not at all) to 5 (extremely) based on how they felt about each one.
3.3 Users’ Overall Experience
To address RQ1 (i.e., What are users’ overall experiences of using EFTeacher?), the usability of the chatbot was assessed using 10 standard SUS items on a 5-point Likert scale. Figure 2 illustrates the distribution of SUS scores. The average SUS average score is 89.93. with a standard deviation of approximately 10.01, indicating that participants generally provided high and consistent usability ratings. Previous research has examined the interpretation of SUS scores in usability assessments, highlighting that higher scores are associated with superior usability experiences [62]. Bangor et al. suggest that products with scores above 70 are generally considered usable, while those in the high 70s to upper 80s indicate better usability. Truly superior products tend to score above 90, reflecting strong user satisfaction. The chatbot’s score of 89.93 places it well within this range, suggesting a highly positive user experience.
Figure 2: SUS scores
3.4 Assessment Regarding Key Content Characteristics
To address RQ2 (i.e., What are users impressions of using EFTeacher?), we focused on four key characteristics of the content of the cue texts. Figure 3 displays four violin plots representing ratings over different time frames (1 month, 3 months, 6 months, 1 year, 3 years, 5 years, and 10 years) for the characteristics of Vividness, Liking, Excitement, and Importance. Each plot shows the distribution of ratings on a scale from 1 to 5, with the width of each violin indicating the density of responses at each rating level.
Ratings for vividness are consistently high across all time frames, with most responses clustered around 4 to 5. This pattern indicates a strong and enduring perception of vividness, with a median rating close to 5 across all intervals.
Similarly, ratings for liking  [^1] How much the participant likes the event show high consistency, predominantly centered near 5 across all time frames. This distribution suggests a lasting sense of appreciation or a positive preference, with little variation over time.
The excitement plot exhibits slightly more variability than was in evidence for vividness and liking. For shorter time frames, such as 1 month and 3 months, ratings are more dispersed, ranging between 3 and 5. However, as time progresses, ratings consolidate towards higher values, with most responses centered around 4 to 5 at longer intervals (e.g., 10 years), indicating increasing consistency and sustained excitement over time.
The importance plot shows a slightly wider spread of ratings for shorter time frames, particularly for 1 month and 3 months, where responses are more varied. Over longer durations, ratings converge near 4 to 5, reflecting a growing or stable recognition of importance with time.
The scores users assigned to cue texts regarding the four key content characteristics highlights the consistently high median ratings for vividness, liking, excitement, and importance across all timeframes, demonstrating that users perceive EFTeacher as impactful based on what it generated for them through their interactions with it.
Figure 3: Ratings for vividness, liking, excitement, and importance characteristics.
3.5 Assessment Regarding User Perceptions
To address RQ3 (i.e., How do users perceive working with an interactive AI chatbot for EFT generation?), the research team employed thematic analysis, i.e., a systematic approach to identify themes from user feedback [32, 63, 64]. Three researchers, each with different areas of expertise, independently reviewed the participants’ open-ended responses multiple times. They each developed a list of codes, which are short labels summarizing key points from the feedback. A facilitator, who had expertise in qualitative methods but limited prior knowledge of the survey (to reduce bias), gathered these codes and compiled them into a comprehensive codebook.
The facilitator and coders then held collaborative sessions to refine the codes, merge similar ones, and identify recurring patterns. This iterative process led to the formation of broader thematic categories. For each theme, the team defined a description and selected representative quotes from user responses to illustrate it. Ultimately, five key categories emerged, which were discussed extensively and finalized in a cohesive codebook.
Using this finalized codebook, the coders reviewed the user responses again, selecting excerpts that best exemplified each theme. The facilitator aggregated these responses and calculated the reliability between coders, a measure of how consistently the coders applied the codes, which was found to be very high, i.e., 94%. The team then agreed on a final table of themes, including example quotes to support each one.
The analysis provided insights into participants’ experiences with EFTeacher, focusing on: communication, adaptability, imagination, reflection, and its interface. These findings highlighted varying levels of engagement and interaction, offering valuable feedback for understanding how users perceived and interacted with the chatbot. The next five subsections explain these themes, providing a framework for understanding participant experiences and interactions with EFTeacher throughout the study. To provide details, quotes from participants are included, with the participant number given for attribution purposes.
3.5.1 Communication
This theme focuses on descriptions of dialogue and communication with EFTeacher. The participants decribe how conversational or human-like their interactions with EFTeacher felt. They widely praised EFTeacher for its ability to engage in human-like conversations. Its tone was frequently described as natural, and exceeding user expectations. Many noted that the dialogue felt impressively lifelike and engaging:
“Actually, the response was better than I would expect a human to respond.” [P105]
“The words that EFTeacher used were just everyday words, nothing fancy or unnatural. This didn’t sound like it was coming from a ’machine’; to me, it sounded natural. I felt as if I was having a conversation with an instructor!” [P112]
Participants emphasized the model’s ability to maintain a conversational flow that mirrored human interactions, often feeling natural [P139], genuine [P137], and incorporating leading questions [P117]. The responses were also praised for being empathetic and relevant:
“It really interacted with me like a human! I expressed so much joy at something it wrote and said I might need to save it to read later, and it responded so thoughtfully.” [P118]
“Frankly, given the same questions (somewhat rote), a smart human being with the same objective as the AI obviously had would come up with similar text.” [P103]
Despite its strengths, some participants found the interactions occasionally too polished or scripted, detracting from the perceived naturalness:
“I think the responses are arguably too well-written to reflect the way the average person speaks. I liked the way it worked a lot, but there are some aspects that don’t feel natural.” [P106]
“Even though it was formed from my prompts, it still sounded very scripted.” [P104]
Repetition also emerged as a common critique. While many users appreciated the flow of the conversation, some noted that EFTeacher tended to repeat similar requests for details, making the interaction feel less dynamic:
“Not only was their storytelling great, but it seemed to flow somewhat like a real conversation, except for the repetitiveness of asking for the details it needed each time.” [P123]
“It was very repetitious. I’m not sure a human would respond the exact same way to every question.” [P119]
These insights suggest that while EFTeacher demonstrates a remarkable ability to create engaging and human-like conversations, there is room for improvement. Participants’ feedback highlights a dual expectation: While many users appreciated EFTeacher’s thoughtful and engaging dialogue, the repetition of similar requests for details caused frustration for some. Striking a balance between maintaining high-quality responses and reducing repetitive patterns will be crucial for future iterative refinement of EFTeacher.
3.5.2 Adaptability, including personalization, customization, and flexibility
Participants widely praised EFTeacher for its ability to respond to personalized prompts and tailor outputs to individual preferences. “I like you can change the cue because if there’s more you want to add or take out you can.” [P122]
This flexibility gave users a sense of control over their interactions and enhanced their overall experience. Many found it easy to adjust the tool’s outputs to align with their specific needs:
“It was very easy to have EFTeacher make changes. I didn’t know if I could ask for specific words to be changed, so I just tried it, and it worked! When I asked for less vivid wording, it changed it to something that I liked better.” [P112]
“I enjoyed how individual words could be replaced. I didn’t like an adjective that was used and asked it to replace that specific word. I also liked how the response could be completely rewritten. It’s great that you can scrap a response entirely and have the bot create a new one.” [P116]
Participant [P118] appreciated how EFTeacher adapted to feedback, allowing them to refine outputs in real time. Simple edits such as rearranging phrases or adjusting tone were implemented with ease:
“The word choice it used was kind of strange, (in one particular instance it said that a table was “laden” with food, and it felt very Middle Ages). I brought this up, however, and EFTeacher easily fixed it! I realize that EFTeacher is able to use the full dictionary to describe things - but that doesn’t always mean the words it chooses are commonly used in our modern vernacular. Again, easy to fix though!” [P118]
Participant [P115] highlighted EFTeacher’s ability to paraphrase and confirm their intent, creating a more engaging and collaborative interaction:
“I liked the way it confirmed what you were trying to convey. It was great that EFTeacher understood my words and paraphrased it back to me.” [P115]
However, one of the participants identified limitations in EFTeacher’s ability to fully simplify responses when requested. While the tool was adaptable, certain adjustments fell short of expectations:
“Even when asked to simplify it was still more than I thought a simple description would be. For some people less words work better for their understanding. So when asked to simplify something, I would expect a shorter answer with right-to-the-point statements.” [P125]
Another participant noted EFTeacher’s capacity to enhance narratives without losing the original intent:
“It picked up on me accidentally submitting an answer before finishing my thoughts. I accidentally hit enter instead of deleting a word and it picked up on that and prompted me to continue my thought” [P122]
Most of the feedback underscores the importance of personalization and adaptability in enhancing user interaction, emphasizing the need to meet diverse user expectations.
3.5.3 Imagination
Participants discussed EFTeacher’s facilitation of imagination, e.g., ability to guide imagery and visualization of future events, describing it as enhancing their milestones and making them feel even more meaningful [P120], painting a vivid picture [P125], and triggering their imagination to visualize these moments [P122]. One participant emphasized its impact:
“I was really struggling to imagine events ten years from now, and it helped me string together some of my piecemeal thoughts!” [P118]
Many users praised its descriptive capabilities, noting that the tool effectively transformed abstract ideas into detailed visuals.
“I enjoyed the details added. I am not a creative writing person and it was nice to see things more vividly.” [P109]
“I liked that the extra details provided a larger picture.” [P138]
“It added vivid details that made it much easier to imagine my future” [P106]
“It was very well done, and I was impressed by the small details it would come up with, and the vividness of its language.” [P123]
Participants appreciated how EFTeacher’s thoughtful use of adjectives and descriptive language made their milestones feel more meaningful and polished:
“Used lots of adjectives. The statement by the AI describes things better than I could put it. It helped me imagine actually being there. By using more descriptive wording I felt as if I could really be doing these things and feeling that way.” [P125]
“The flowery words used helped see how great the future may be.” [P115]
Several participants highlighted that the guided prompts and questions posed by EFTeacher were particularly effective in facilitating the imagination of their stories. The questions were described as helpful [P109], well-designed [P103], thought-provoking [P119], and structured yet flexible, allowing for creativity:
“The program/EFTeacher asked perfect questions to help me envision these events!” [P137]
“I liked how it asked questions about my storyline because it made me think about the setting, people, and everything. It really helped bring out the writer in me.” [P122]
Participants also found the tool easy to use, noting that it naturally guided them into crafting imaginative stories. As one participant described:
“It was easy to use, and coax’s it into imaginative stories. It was just very simple, as long as you understood what it was asking for, to use and make it come up with very nice stories.” [P123]
One participant mentioned expecting the chatbot to take the lead in storytelling or provide meaningful new ideas beyond their initial input:
“It felt like I was the one coming up with every story, and they were just riffing on what I came up with by myself. I don’t think this is bad, but I didn’t feel like it was giving me anything to go off of besides the timeframe the story took place in.” [P123]
EFTeacher’s ability to stimulate the imagination through vivid descriptions, questions, and prompts is a standout feature. It has the potential to further enhance its role as a facilitator of imagination and creativity, making it an even more effective tool for users seeking to envision and articulate their future.
3.5.4 Reflection
Some participants, reflecting on their experience, expressed emotions, connections, and results of the imagery elicited by EFTeacher. They shared their key takeaways from interacting with EFTeacher, highlighting how it prompted them to reflect on their goals and aspirations in new ways.
One participant reflected on how EFTeacher helped them think more broadly about their future:
“I don’t typically think very far into the future, and it made me consider where I want my life to be in X months/years.” [P106]
Similarly, another participant noted how EFTeacher prompted thoughts they had not previously explored:
“It made me think about things I really haven’t thought about.” [P122]
“The system made realize what the future may hold.” [P115]
The reflective process elicited by EFTeacher was described as both emotionally impactful and validating. For instance, one participant shared how the experience left them feeling hopeful:
“It made me feel hopeful about these future events. It was also validating some of the feelings that I had.” [P107]
This sentiment was echoed by another participant, who highlighted the excitement generated by envisioning future events:
“ it made you excited for the event.” [P138]
[P107]’s feedback highlighted the vividness of the experience, which played a key role in inspiring hope:
“It was so vivid that it gave me hope.” [P107]
“The AI was on point and gave me a lot of hope that I went to work the next day and could imagine these events” [P107]
EFTeacher’s interactive and specific prompts encouraged self-reflection by helping participants gain meaningful insights from their experiences. Participants mentioned how this approach encouraged self-reflection:
“Being prompted to imagine specific details made me reflect on my life.” [P137]
“It allowed me to express my dreams of the future.” [P127]
“The way the system took my thoughts and made them into a great story. The use of descriptive words brought my true-to-heart feelings to life with little effort on my part.” [P115]
Another participant [P118] reflected on their overall experience with EFTeacher:
“I expressed so much joy at something it wrote… and it responded thoughtfully.” [P118]
One participant appreciated the experience but found that redundancy slightly reduced their enjoyment:
“I really liked this experience. The redundancy made sense, but it was the only aspect that made it feel slightly less enjoyable for me. That said, I understand that not everyone envisions their personal journey in the same way or for the same duration.” [P106]
For another participant, the level of detail felt overwhelming at times:
“Some of the responses contained so much detail that it became overwhelming. There was just a lot of information to process.” [P138]
These observations illustrate how EFTeacher could guide participants to envision their futures with greater clarity. By prompting users to focus on specific details and engaging with their inputs thoughtfully, EFTeacher inspired reflections and a connection to their aspirations.
3.5.5 Interface
Participants shared some insights on EFTeacher’s interface, including its usability and visual design, discussing both its strengths and areas for improvement. While many found the platform intuitive and easy to navigate, others highlighted specific challenges that impacted usability and accessibility.
One common suggestion was to improve the design for better readability, and suggestions for better color contrast were common. For example, participant [P129] commented on the color scheme and suggested improved audio as well as visual support:
“The color scheme needed improvement, as some words blended into the background, making them hard to see.” [P129]
“no option for audio and visual. Hearing the descriptions and scenarios aloud would have added good detail” [P129]
Beyond readability, participants proposed features to enhance the interface’s usability. For instance, the inclusion of a word count feature was suggested to help users meet expectations for detail:
“If there is an expectation of level of detail or goal, then it would be nice to track that from the start.” [P117]
Another participant expressed difficulty with copying text, emphasizing the need for smoother interaction:
“I didn’t like how hard it was to copy what was said. I had a hard time copying the rewrites.” [P137]
Voice control was also suggested as a potential accessibility feature to simplify interactions:
“Voice control. Make it easier.” [P137]
Beside these, participants found it easy to use [P119], [P120], [P136]. One participant found the interface intuitive and appreciated the straightforward instructions:
“It was easy to input information, and the instructions were clear.” [P124]
While the interface was appreciated for its ease of use, incorporating participant feedback could further enhance the user experience across diverse individuals.
4 Discussion
The goal of this study is to explore users’ experiences and perceptions of EFTeacher, an AI chatbot designed for EFT generation. Specifically, we aim to understand users’ overall experiences with the chatbot, their impressions of its functionality, and their perceptions of working with an interactive AI tool for EFT generation.
Our analysis indicates that the chatbot’s descriptive language was a polarizing feature among participants. Some individuals appreciated its vivid descriptions, which enhanced their sense of vividness and made their interactions feel richer. Participants highlighted the chatbot’s ability to create detailed visuals and make scenarios feel more tangible. However, others felt that the language was overly elaborate and occasionally unnecessary. For these users, the level of detail made conversations less natural.
Interestingly, this issue could potentially be mitigated through the chatbot’s adaptability features, which allow users to request simpler or less descriptive responses. However, it became evident that participants did not utilize it frequently, suggesting a gap in feature accessibility or motivation to adjust responses mid-interaction. The chatbot’s adaptability was highlighted as a key strength by some partiicpants, allowing them to refine responses and tailor interactions to their preferences. However, some participants did not utilize these features effectively, leading to missed opportunities for improving their experiences. Multiple engagements with the chatbot could facilitate a learning process that helps users become more familiar with these adaptability features, ultimately improving their ability to personalize interactions.
Participants frequently highlighted the positive emotional impact of their interactions with the chatbot, emphasizing its ability to inspire hope, optimism, and validation. The chatbot’s generated EFTs encouraged users to reflect on their goals and aspirations, often guiding them to explore aspects of their future they might not have otherwise considered.
The vividness and personalization of the chatbot’s EFTs played a key role in creating these positive emotional experiences. EFTeacher demonstrated potential in helping users articulate their thoughts with greater detail and structure, encouraging them to engage more deeply with their personal visions. This approach not only strengthened their emotional connection to the conversation but also nurtured a sense of optimism about the future.
This aligns with the core principles of EFT interventions, where individuals vividly imagine and describe positive future events. This process suggests that an AI chatbot could be a valuable tool for effective EFT interventions by encouraging users to consider the future and helping them focus on their long-term goals.
While the reflective nature of the interactions was highly valued, some participants noted that the level of detail could occasionally feel overwhelming and wordy. This stems from AI-generated language, which, coming from a machine, does not always sound natural. However, despite these constraints, AI can still be a valuable tool for structuring and refining EFT-related written content, making it a supportive tool, especially for those who have difficulty with writing. Participants mentioned that the repetition in the chatbot’s responses negatively impacted their interactions. This repetition stemmed from the chatbot prompting users for seven different events at seven different times, following the same process each time. Participants described this as overly repetitive and excessive. Reducing repetitive patterns could improve the chatbot’s ability to deliver more natural, human-like interactions. To improve user experience, our research group has decided to conduct the EFT teacher interview across four (instead of seven) different time frames in future iterations.
5 Design Insights
5.1 Integrating Voice Capabilities into EFTeacher
Based on the findings from our thematic analysis, integrating voice capabilities into EFTeacher could enhance its effectiveness in creating a more engaging and natural conversational experience. Participants consistently praised EFTeacher’s ability to emulate human-like conversation, with its tone being described as thoughtful and lifelike. Introducing voice as a medium of interaction could address these limitations by enhancing the emotional and contextual depth of communication.
Prior research has supported the importance of voice in human-agent interaction. According to [65], voice impacts not only the content of the speech but also the overall perception of the agent’s personality and context-specific relevance. Additionally, audio delivery has been explored in the context of Episodic Future Thinking (EFT) interventions with promising results. Studies such as [25, 17, 66] support the use of audio in EFT interventions. Participants can generate cues and then vividly imagine or listen to audio recordings of their EFT cues during tasks that assess delay discounting. This approach has demonstrated effectiveness in enhancing the intervention experience by leveraging audio’s ability to evoke richer emotional and cognitive engagement.
A scoping review of the literature further suggests that voice assistants can serve as a valuable tool in the research of behavioral health interventions [67]. This supports the potential for EFTeacher to leverage voice capabilities not only to enhance user engagement but also to support interventions aimed at improving behavioral and emotional well-being.
Voice integration also addresses accessibility concerns raised during usability testing. For example, participants suggested features like voice control to simplify interactions, especially for users who might find typing or reading responses challenging. Voice output and input would not only improve usability but also expand EFTeacher’s accessibility to a broader audience.
5.2 Leveraging Visual Capabilities for EFTeacher
Advancements in LLMs present a promising avenue for building effective human-AI interaction. The use of multimodal features of LLMs is increasingly gaining attention in recent technologies, as demonstrated in [68, 69, 70]. One such feature is the ability of models to generate images from text prompts, as explored in [69]. Integrating this capability into EFTeacher could allow the chatbot to produce visual EFTs in response to user queries, making interactions more engaging and particularly useful for tasks that require visual representations, such as illustrating future events based on user prompts and generating written EFTs.
The benefits of incorporating visual elements are further supported by research on EFT. A study [71] investigated the effects of drawn and written future cues, and found that both drawn and written future cues led to reduced delay discounting compared to the control group, with drawn cues offering unique advantages in terms of personal relevance and the vivid encoding of details. These findings suggest that enabling EFTeacher to generate or support visual representations of future scenarios could further enhance its impact on future-oriented decision-making.
Moreover, incorporating visual elements could address a key finding from our study: while some participants appreciated the AI-generated language, others found it overly elaborate or unnecessary. This discrepancy highlights the need for alternative modes of interaction, such as visualizations, for users who may not connect with the AI chatbot’s language or who prefer a less verbose experience. By offering visual representations alongside text-based descriptions, EFTeacher could accommodate a broader range of user preferences.
6 Limitations and Future work
For this study, we asked participants to describe seven future events. This approach helped us gather more AI-generated EFT data, but made the experience feel repetitive for some participants. In future studies, we plan to reduce this to four events, which we believe will make the interaction more positive and less repetitive while still collecting valuable data.
Our study design only measured immediate user perceptions and experiences. We did not track longer-term behavioral or psychological outcomes. In subsequent research, the research team aims to conduct longer-term studies to observe how continued use of EFTeacher’s generated cues might influence users’ behavior over time. This longitudinal approach will allow us to assess whether the chatbot’s interactive EFT features have a sustained impact on users’ motivation, goal-setting, and health-related behaviors.
In parallel, future research should compare user perceptions and the efficacy of EFTeacher to the existing survey method and clinician-led interviews. Although the survey method was designed as a more scalable alternative to the human interview, poor cue quality cues in this self-administered method (e.g., vague or non-episodic) may reduce treatment fidelity and thus efficacy of the intervention. Although EFT is a promising behavioral intervention to facilitate weight loss and dietary [17, 18, 19, 20, 21], significant effects on health behaviors are not always observed [20, 19]. Thus, future research should examine whether the method used to generate cues moderates these behavioral effects and user perceptions.
Despite its limitations, our study offers initial evidence supporting the feasibility and promise of a chatbot-based EFT intervention, suitable to guide follow-on studies.
7 Conclusion
In this study, we evaluate EFTeacher, an AI chatbot built on GPT-4-Turbo, designed to assist participants in identifying meaningful future events and creating detailed text descriptions (cues) for those events. We investigated users’ overall experiences with EFTeacher, their impressions of the chatbot, and their perceptions of working with an interactive AI tool for EFT generation.
Our findings highlighted both strengths and areas for improvement in EFTeacher’s design and functionality. Quantitative analysis of user impressions, represented through content charectristics of vividness, liking, excitement, and importance ratings, revealed positive trends.
Analysis of user feedback revealed five key themes that provided deeper insights into users’ experiences with EFTeacher. Participants appreciated the chatbot’s vivid and descriptive language, which enhanced engagement and encouraged meaningful self-reflection. This feature helped users articulate their goals and vividly visualize future scenarios. However, some participants found the detailed language excessive, emphasizing the need for adaptability in response style. The chatbot’s adaptability features—enabling refinements in tone and style—were identified as a significant strength. However, these features were underutilized by some participants, indicating the need for better user guidance. Participants also highlighted challenges with repetitive task structures. Reducing repetitive patterns could improve usability and promote a more natural conversational experience.
Overall, EFTeacher demonstrated potential as an effective tool for guiding users through EFT generation. Its responses were highly valued, with participants emphasizing its ability to inspire reflection on their goals. The study identified areas for improvement, such as minimizing repetitive patterns and offering clearer guidance on adaptability features. Our design suggestions could further enhance the application of AI chatbots based on LLMs in EFT interventions for maladaptive health behaviors.
Acknowledgment
This work was funded by the Seale Innovation Fund and the National Institute of Diabetes and Digestive and Kidney Diseases (5R01DK129567-02S1).